ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 113, 1999

Alesker, S. On the measure of intersection of cylinders  231

Barros, M., J. L. Cabrerizo and M. Fernández Reduction of variables for Willmore–Chen submanifolds in seven spheres

29

Brendle, J. Between P-points and nowhere dense ultrafilters  205

Cabrerizo, J. L. See M. Barros

Cholak, P. A. and A. Nies Atomless r-maximal sets  305

Džamonja, M. and S. Shelah ♣ does not imply the existence of a Suslin tree

163

Fernández, M. See M. Barros

García-Sánchez, P. A. See J. C. Rosales

Gross, B. H. Algebraic modular forms  61
Loke, H. Y. Dual pairs correspondences of E_{8,4} and E_{7,4}  125
Mao, Z. and S. Rallis On a cubic lifting  95

Matouš ek, J. and P. Valtr On visibility and covering by convex sets

341

Neumann, K. A criterion for the primeness of ideals generated by polynomials with separated variables

1

Nies, A. See P. A. Cholak

Park, K. K. On directional entropy functions  243

Petrogradsky, V. M. Exponential generating functions and complexity of Lie varieties

323

Rallis, S. See Z. Mao

Randrianantoanina, B. On isometric stability of complemented subspaces of L_{p}

45

Regev, A. Asymptotics of degrees of some S_{n}-sub regular representations

15

Rosales, J. C. and P. A. García-Sánchez Presentations for subsemigroups of finitely generated commutative semigroups

269

Shelah, S. See M. Džamonja

Siburg, K. F. Aubry–Mather theory and the inverse spectral problem for planar convex domains 285 Valtr, P. See J. Matoušek